OpenStudy (anonymous):

What is the simplified form of x-3/x^2+x-12 multiply x+4/x^2+8x+16

OpenStudy (anonymous):

\[\frac{ x-3 }{(x-3)(x+4) } * \frac{ x+4 }{(x+4)^{2} }\] after simpilyfying this will be equal to \[\frac{ 1 }{ (x+4)^{2} }\]

OpenStudy (anonymous):

you know how did I factorize the equations right ?

OpenStudy (anonymous):

@Avva eh not really

OpenStudy (campbell_st):

@Avva, no need to know how to factorise... its already done...

OpenStudy (anonymous):

@campbell_st so the answer is 1/(x+4)^2

OpenStudy (campbell_st):

I'll let you calculate it... by factoring the quadratics

OpenStudy (anonymous):

well in this problem you can multiply the numerator and it will be simplified with the denominator first bracket leaving x^2+8x+16 but you might need factorization for solving other problems to solve any quadratic equation \[a x^{2} + b x +c =0\]you apply in this formula \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{2a }\] you may see non-factorable equations (complex numbers it is OK if you haven't studied it yet) it will take you some time so you don't have to apply in the formula if you are given a simple equation try to figure it out without it e.g. \[x ^{2}-x-12 = 0\] so you ask yourself what are the two numbers say (m,n) m*n = -12 m+n=-1 on practice you will be great at that sorry for talking a lot I 'll google a better way to solve quadratic equations and tell you :)